### v1

# Markovian Quantum Computation

Abstract
We propose a general prototype of quantum computation based on the finding of many-body ground states. The scheme is inspired by cavity cooling, involving the emulation of a zero-temperature reservoir. Repeated discarding of ancilla reservoir qubits extracts the entropy of the system, driving the system towards its ground state. At the same time, measurement of the discarded qubits hints at the energy level structure of many-body systems. We show that quantum computation based on this Markovian process is equivalent in its computational power to the one based on quantum circuits. We exemplify the scheme with several illustrative use cases for combinatorial optimization problems, discussing the efficiency and the issue of local energy minima. We also discuss its application to the preparation of quantum many-body ground states for gapped systems.
I Introduction
Computing many-body ground states is extremely hard in general: for classical Hamiltonians, it falls into the NP-hard complexity class, and for quantum Hamiltonians, it is classified as QMA-hard [1, 2, 3]. The difficulty arises mainly from the geometric frustration, i.e., the stark mismatch between local and global energy minima. The problem is combinatorial in nature: one needs to find the optimal configurations of data among exponentially many, mutually conflicting ones. This is indeed the core challenge of combinatorial optimization problems [4]. It is well-established that combinatorial optimization problems can be reduced to the task of finding the ground states of Ising-like spin models [5, 6, 7].
It is widely believed that quantum computers are not universally efficient in solving general ground-state problems, even for classical Hamiltonians. Nonetheless, we still need to find out specific classes of such problems that benefit from quantum computation and are yet practically useful. Also, there’s a need to devise diverse quantum algorithms for them, each having its unique advantages. We have a number of tools available for such purposes [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] (Ref. [21] contains a nice summary of earlier schemes). The most prominent is adiabatic quantum computation (AQC) [10, 11, 12]. The AQC exploits the adiabatic evolution of the instantaneous ground state when the Hamiltonian is adiabatically varied. The time scale of the operation is thus governed by the minimal spectral gap during the evolution. However, analyzing the spectral gap is extremely difficult, leaving the performance of AQC largely unknown. There are also hybrid approaches, such as quantum approximate optimization algorithms (QAOA) [14, 15, 13] and variational quantum eigensolvers (VQE) [16, 17, 18], aided by classical optimization subroutines.
As mentioned, the difficulty in analyzing AQC arises as the initial and final Hamiltonians intermingle. On the other hand, the properties of the final Hamiltonian alone can be largely unveiled, even though its ground state remains unknown. Given this, it is natural to consider schemes to directly cool the final Hamiltonian system to its ground state. While earlier schemes based on dissipative engineering [19, 20, 21] appear to achieve this goal, a closer inspection reveals that they actually lack desirable properties as an ideal cooling method. To clarify this, it is instructive to discuss classical and quantum Hamiltonians separately. By the former, we mean those Hamiltonians diagonalized in the computational basis.
For classical Hamiltonians, the earlier dissipative schemes do not provide any quantum advantage. There are a few fundamental reasons. They implement a completely-positive trace-preserving (CPTP) map that transforms the state $\rho$ to $\sum_{k}E_{k}\rho E_{k}^{\dagger}$ [22]. Here, each Kraus operator $E_{k}$ directly accesses the associated local part of the system, driving it into a certain direction. However, as the energy of the whole system is a global quantity, determining the optimal direction based solely on a small part is not possible (note that this was the very origin of the geometric frustration, rendering the problem intractable). For this reason, the state is randomized when the measurement projects the state into an undesired subspace. Note, however, that this entire process is essentially equivalent to a mere random guess. Furthermore, for classical Hamiltonians, they do not involve any operations in non-computational bases. This means that they are, in fact, identical to classical probabilistic models. In the case of quantum Hamiltonians, the problem is more involving: one need to find an appropriate basis containing the ground state. This is the place where the above CPTP map comes into play. Apart from the role of fixing the basis, however, the above mentioned limitations persist.
An optimal cooling should involve quantum transitions that take place exclusively in the direction of lowering energy. Here, a mechanism to prevent heating can be ensured only by an external agent, which globally accesses the system, gathers its energy, and relaxes it into a zero-temperature Markovian reservoir. This process is exactly given by cavity cooling methods [23]. The aim of this paper is to materialize this concept as a general quantum computation model and discuss its efficiency in some illustrative scenarios. Our model can be paralleled with AQC for their similarity in motivation. On the other hand, unlike AQC that is a unitary process, our model exploits a Markovian process. In order to stress this and for the sake of brevity, we will refer our model to Markovian quantum computation (MQC) hereafter.
Being a Markovian process, MQC inherits the intrinsic advantages of the earlier dissipative schemes, alleviating the demand of precise system controls to some extent. On top of this, MQC takes true quantum advantages. In fact, we show that MQC is equivalent to quantum circuit models in terms of computational power. This argument is similar to the established one regarding the computational equivalence between AQC and quantum circuit models [11]. All three frameworks are thus computationally equivalent. However, this assertion needs to be interpreted with care. For instance, the equivalence of AQC to MQC is established by translating the unitary evolution of AQC into a quantum circuit, which is then converted into a Hamiltonian for MQC. However, this resulting Hamiltonian differs from the original problem Hamiltonian of the AQC. It is unclear if the feasibility of ground-state preparation by AQC implies that of the same Hamiltonian by MQC, and vice versa.
Any ground-state preparation scheme, including AQC, inevitably suffers from the issue of local energy minima [6, 24, 25]. In MQC, one can design tunneling transitions out of local minima if the information is given on their nature. This flexibility also constitutes the advantage of MQC. In practice, one could try various transition terms in the Hamiltonian, while observing the occurrence of the transition from the detection of cavity photons. This, in turn, provides a method to inspect the energy-level structure of the Hamiltonian.
In this work, our main focus will be placed on MQC for combinatorial optimization problems. Such problems involve classical Hamiltonians and have a number of favorable features for MQC. We consider two extreme cases. The first case is free from local energy minima, but the transition rate is superpolynomially small in the system size. This case turns out to be equivalent to Grover’s quantum search algorithm. The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begins to pose challenges. In this case, high-order transitions, required to get out of local energy minima, become the bottleneck. These transitions require a time that is exponentially large in the order, which can increase with the system size in general. Finally, we briefly discuss the case of quantum Hamiltonians. This case is heavily system-dependent and harder to analyze. We argue that the spectral gap above the ground state determines the basic time scale of the cooling.
II Notation and working principles
Consider an $N$-qubit system described by Hamiltonian
$H_{P}=\sum_{z=0}^{2^{N}-1}E(z)|z\rangle_{P}\langle z|,$ (1)
where $E(z)\leq E(z^{\prime})$ for $z<z^{\prime}$. Let us call $H_{P}$ a problem Hamiltonian. Our aim is to find the ground state $|0\rangle_{P}$. Transitions among the energy levels $|z\rangle_{P}$ are allowed by introducing in the Hamiltonian what we call a transition term $H_{T}$, which does not commute with $H_{P}$. $H_{T}$ should be designed to suit the need of the given problem. We will consider two examples below. Here, we assume the energy level spacings of $H_{P}$ are characterized by an energy scale $\Delta>0$, and so is the norm $\|H_{T}\|$. Our aim is to introduce a parameter $0<\lambda\ll\Delta$ to treat $\lambda H_{T}$ as a perturbation to $H_{P}$.
If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle$ to $|z^{\prime}\rangle$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|\langle z^{\prime}|\lambda H_{T}|z\rangle|$, ignoring the influence of the degeneracy. In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|\langle z^{\prime}|\lambda H_{T}|z\rangle|$ and couple it to the transition. Let us denote by $|\cdot\rangle_{m}$ the photon number state of the $m$-th cavity. If the cavity is initially empty, the transition from $|z\rangle_{P}|0\rangle_{m}$ to $|z^{\prime}\rangle_{P}|1\rangle_{m}$ is allowed only when $E(z^{\prime})\simeq E(z)-\omega_{m}$, providing a mechanism to avoid heating.
Introducing $M$ different cavity modes, our final Hamiltonian reads
$\begin{split}H&=H_{P}+\sum_{m=1}^{M}\omega_{m}c_{m}^{\dagger}c_{m}\\ &\quad+\lambda H_{T}\otimes\left[a_{0}I+\sum_{m=1}^{M}(c_{m}+c_{m}^{\dagger})\right].\end{split}$ (2)
Here, $a_{0}\in\{0,1\}$ determines whether the transitions preserving the cavity photon numbers are allowed. This term helps the state escaping from local minima through high-order transitions. In practice, it is generally enough to restrict the maximum number of photons in each cavity to one. In this case, $M$ cavity modes can be simulated with $M$ qubits.
Initially, the system is prepared in state $2^{-N/2}\sum_{z=0}^{2^{N}-1}|z\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$. A single cycle of cooling is performed by unitarily evolving the system according to Hamiltonian (2) for an appropriate duration of time, measuring the cavity states, and resetting every cavity to the vacuum state. Here, the measurement of the cavity photons is not mandatory, but it would reveal some information on the current state and the internal transition structure of the problem Hamiltonian.
The computation is done by repeating this cycle. The required number of cycles depends on the problem and strategy. One possible strategy is to repeat until the cavity photon is not detected for an enough number of cycles. This signals with high probability that the system has reached either the ground state or local minima. For NP problems that are guaranteed to have at least one solution, deciding which is the case is straightforward. Otherwise, one may repeat the same computation to increase the confidence. Given the unknown ground state not being verifiable in general, the decision is probabilistic after all.
III Equivalence to circuit models
Figure 1: The transition diagram into which a quantum circuit is converted for MQC. $|\phi_{t}\rangle_{P}$, $|t\rangle_{C}$, and $|\cdot\rangle_{1}$ denotes, respectively, the state of the quantum circuit at the $t$-th time step, the clock state to label the time step, and the number state of the cavity. $\Delta$ is the energy difference between the adjacent levels and $\lambda$ is the Rabi frequency of the designated transition. The wavy lines represent the transition after detecting and clearing the cavity photon.
Ref. [11] shows that AQC is equivalent to the quantum circuit model in computational power up to polynomial overhead. The forward direction of the equivalence is rather straightforward because the entire procedure of AQC is a unitary transformation, which can be simulated efficiently in quantum circuits. Showing the other direction replies on the clock state $|t\rangle_{C}$ that is tagged to the state at the $t$-th time step of the quantum-circuit computation.
A similar equivalence relation can be shown for MQC. Again, showing that MQC can be simulated efficiently in quantum circuits is straightforward. We thus focus on the other direction. Suppose a certain quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step. Let $|\phi_{t}\rangle_{P}\equiv U_{t}^{P}U_{t-1}^{P}\cdots U_{1}^{P}|\phi_{0}\rangle_{P}$ be the state after applying $U_{t}^{P}$, and tag this state with the clock state as $|\phi_{t}\rangle_{P}|t\rangle_{C}$. The idea is to get $|\phi_{t}\rangle_{P}|t\rangle_{C}$ transformed into $|\phi_{t+1}\rangle_{P}|t+1\rangle_{C}$ in the course of an energy-lowering transition. This is enabled by choosing the terms in Hamiltonian (2) as
$H_{P}$ $=-\sum_{t=0}^{T}t\Delta|t\rangle_{C}\langle t|,$ (3)
[…]
where the parameter $\Delta>0$ is an arbitrary energy scale, and the other parmeters are chosen as $M=1$, $\omega_{1}=\Delta$, and $\alpha_{0}=0$. If the initial state is chosen to be $|\phi_{0}\rangle_{P}|0\rangle_{C}|0\rangle_{1}$, the entire transition dynamics is restricted to a subspace spanned by $\{|\phi_{t}\rangle_{P}|t\rangle_{C},0\leq t\leq T\}$, as depicted in Fig. 1. It can be seen that the average time needed to reach the ground state, which is the final state of the quantum circuit, is polynomial in $T$. As the transition rate $\lambda$ is restricted by the energy gap $\Delta$, the computation time is $\mathcal{O}(T/\Delta)$.
IV Combinatorial optimization by Markovian quantum computation
IV.1 Notation
Consider combinatorial optimization problems that aim to find out the optimal solutions from a set of $2^{N}$ possible configurations represented by $N$-bit string $z\equiv z_{1}z_{2}\cdots z_{N}$. The optimality of each configuration is evaluated by a non-negative cost function $E(z)$. The objective is to identify configurations $z$ that minimize $E(z)$. This problem is straightforwardly translated into the ground-state finding problem, where the cost function $E(z)$ is identified with the energy eigenvalue in Hamiltonian (1). Here, we set $E(z)$ to be integer multiples of a parameter $\Delta>0$. The cavity frequencies in Hamiltonian (2) can then be chosen as integer multiples of $\Delta$. By choosing $M$ to be the maximum of $E(z)$, all possible transitions in the problem Hamiltonian can be coupled to a cavity transition. This imposes a polynomial overhead to the computational resource.
IV.2 Reproduction of Grover’s search algorithm
Grover’s quantum search algorithm is paradigmatic in quantum information theory [26, 22]. It can be recast as an algorithm to find the zeros of a binary function $f(z)\in\{0,1\}$ from the set of $2^{N}$ configurations $z$. Provided that $f(z)$ is computed only by an external agent, called the oracle, Grover’s algorithm requires $\mathcal{O}(2^{N/2})$ oracle calls to find the solution, while the number becomes $\mathcal{O}(2^{N})$ for the best classical algorithm. The quadratic speedup of Grover’s algorithm is proven to be optimal unless the structure of the function $f(z)$ is exploited somehow [27, 28, 29].
To perform an analogous task in MQC, we adjust the Hamiltonian (2) as
$\begin{split}H&=\sum_{z=0}^{2^{N}-1}\Delta f(z)|z\rangle_{P}\langle z|+\Delta c_{1}^{\dagger}c_{1}\\ &\quad+\lambda\bigotimes_{i=1}^{N}\frac{I_{i}+X_{i}}{2}\otimes(c_{1}+c_{1}^{\dagger}),\end{split}$ (5)
where $I_{i}$ and $X_{i}$ denote the identity and Pauli $X$ operators acting on qubit $i$, respectively. Let $n_{0}$ be the number of zeros of $f(z)$. For the initial state $2^{-N/2}\sum_{z}|z\rangle_{P}|0\rangle_{1}$, the problem Hamiltonian $H_{P}$ effectively becomes two-dimensional, where the Hilbert space is spanned by
$|\phi_{0}\rangle_{P}=\frac{1}{\sqrt{n_{0}}}\sum_{f(z)=0}|z\rangle_{P}$ (6)
and
$|\phi_{1}\rangle_{P}=\frac{1}{\sqrt{2^{N}-n_{0}}}\sum_{f(z)\not=0}|z\rangle_{P}.$ (7)
In this basis, the off-diagonal element becomes ${}_{P}\langle\phi_{0}|\lambda\bigotimes_{i}\frac{I_{i}+X_{i}}{2}|\phi_{1}\rangle_{P}\simeq\lambda n_{0}^{1/2}2^{-N/2}$. The computation time is thus $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm. One advantage of MQC implementation is that the end of computation is signaled by a detection of the cavity photon. This is especially beneficial when the number of solutions $n_{0}$ is unknown.
In this example, the problem Hamiltonian has only two energy levels corresponding to the solution states and the rest. A naturally following question is if the performance can be improved when the problem Hamiltonian has more than two energy levels. It can be seen that with the transition term chosen above, more energy levels just make the computation slower. To see this, note that the quadratic speedup comes from the collective transition, the rate of which increases with the degeneracy of the energy level. The increase in the number of energy levels leads to the decrease in the degeneracy of each level, which diminishes the collective effect. Moreover, it results in an increased number of transitions required to reach the ground state.
IV.3 Alternative approach
In the previous example reproducing Grover’s search algorithm, the transition term generates an all-to-all interaction with the identical transition strength. This makes the problem Hamiltonian completely featureless except for each energy level having a different degeneracy. While the absence of any local energy minimum is advantageous, the transition strength decreases superpolynomially with the number of qubits.
In this subsection, we consider an opposite limit, where the transition term is given by
$H_{T}=\sum_{i=1}^{N}X_{i}.$ (8)
The off-diagonal element ${}_{P}\langle z|H_{T}|z^{\prime}\rangle_{P}$ is then non-vanishing and equals one, independently of $N$, only when the Hamming distance between $z$ and $z^{\prime}$ is one. While the transition strength is maximum, local energy minima now become an issue. As the distance between two configurations is at most $N$, any state $|z\rangle_{P}$ can reach the ground state in at most $N$ transitions in principle. However, we do not know a general rule to take the right path for every state. For this reason, we take the direction of lowering the energy, even though it does not coincide with the direction of approaching the ground state. This mismatch is the origin of local energy minima.
Note that every state $|z\rangle_{P}$ is linked to $N$ different states by transition. It is instructive to envisage this as an $N$-regular graph with $2^{N}$ vertices. Each vertex is endows with a potential determined by the problem Hamiltonian, and the transition occurs along the edges toward the direction of not increasing the potential. Once the population is trapped in a local potential minimum, a high-order transition is needed to get out of it.
To elucidate this mechanism, consider an $n$-th order transition through the sequence $|z_{0}\rangle_{P}\rightarrow|z_{1}\rangle_{P}\rightarrow\cdots|z_{n}\rangle_{P}$. For the moment, suppose that $\alpha_{0}=1$ in Hamiltonian (2) and there is no cavity mode involved. For such a transition to occur, the first requirement is $E(z_{0})=E(z_{n})$. The perturbation theory yields the effective transition rate $\mathcal{O}[\lambda(\lambda/\Delta)^{n-1}]$. However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ are different, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition. Consequently, the most prominent transitions come from the cases wherein the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$. Now taking the cavity modes into account, the last condition can also be met when an intermediate transition $|z_{i}\rangle_{P}\rightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\rightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$. This is always possible provided $E(z_{i+1}^{\prime})<E(z_{i+1})$ and the appropriate cavity mode exists. By incorporating multiple cavity modes, various tunnelling-out transition channels can appear. For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$ are expected to play a crucial role in overcoming local minima.
To illustrated the above idea, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [30, 22]). To be specific, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem. The problem Hamiltonian is constructed as follows. Consider a multiplication of two 3-bit integers $\bar{x}\equiv(x_{2}x_{1}x_{0})_{2}$ and $\bar{y}\equiv(y_{2}y_{1}y_{0})_{2}$ resulting in a 6-bit integer $\bar{z}\equiv(z_{5}z_{4}\cdots z_{0})_{2}$, incorporating additional four carry bits $\bar{c}\equiv(c_{3}c_{2}c_{1}c_{0})_{2}$. For a given value of $\bar{z}$, we turn each step of the calculation into an energy term in the problem Hamiltonian. Here, 10 qubits are needed to encode $\bar{x}$, $\bar{y}$, and $\bar{c}$, while $\bar{z}$ is hard-coded in the problem Hamiltonian. For example, the first energy term has value zero if $x_{0}y_{0}=z_{0}$ and $\Delta$ otherwise, the second has value zero if $x_{1}y_{0}+x_{0}y_{1}=(c_{0}z_{1})_{b}$ and $\Delta$ otherwise, and so on. The ground state, representing the solution of the factoring, has energy zero as it satisfies all the conditions, and has a two-fold degeneracy as $\bar{x}$ and $\bar{y}$ are interchangeable.
Figure 2: Numerical simulation of MQC for integer factoring of 35. Dotted curves represent the energy with respect to the number of cooling cycles, averaged over $10^{3}$ samples, for (a) $\alpha_{0}=1$ and (b) $\alpha_{0}=0$. Gray curves represent typical individual trajectories.
Fig. 2 shows the results of our numerical simulation for $\bar{z}=35$. We take the transition term as in Eq. (8) and set the parameters in Hamiltonian (2) as $M=3$, $\omega_{m}=m\Delta$, $\lambda=\Delta/10$, and $a_{0}=1$. The duration of a single cooling cycle is chosen to be $\pi/2\lambda$. The simulation is performed by exact diagonalization. Fig. 2(a) shows that the energy of the system vanishes in time, indicating that the system evolves into the ground state. As a comparison, we have changed $a_{0}$ to zero in Fig. 2(b) to test the effectiveness of the aforementioned mechanism for tunneling out from local minima. In the absence of such a mechanism, the system almost always evolves into local minima, which appear to be present in the first excited level.
A remark is in order. Our choice of the parameter $\lambda=\Delta/10$, chosen to fulfill the requirement that the transition rate is much smaller than the energy difference, appears to be effective, as demonstrated in Fig. 2(a). However, it should be noted that the optimal choice of $\lambda$ depends on the characteristics and size of the problem. In general, smaller $\lambda$ is preferable as the collective effect enhance the transition rate. This effect is particularly important for combinatorial optimization problems, which are characterized by large degeneracies in the problem Hamiltonian. As the degeneracies of low energy levels are much smaller in general, the collective effect is also weaker in the low-energy part of the spectrum. This means that there exists a parametric regime of $\lambda$ where the heating is sufficiently avoided for the ground state, but not for higher energy states. On the other hand, by choosing smaller $\lambda$, leading to slower transitions, heating could be avoided throughout the entire energy spectrum. While the cooling works in both cases, it is unclear which one is more efficient.
V Discussion
In this work, we have introduced the notion of Markovian quantum computation, along with a few illustrative examples. The idea is quite straightforward: cooling the system with a problem Hamiltonian through energy-lowering transitions, as in cavity cooling. An essential part here is to choose an appropriate transition term that generates off-diagonal elements in the diagonal basis of the problem Hamiltonian. There exists a trade-off between the connectivity and strength of the transitions. We have discussed two opposite limits when the ground state is the solution of a combinatorial optimization problem. In both cases, the running time generally increases superpolynomially with the system size. It is unclear if the performance can be substantially improved by interpolating between the two opposites.
MQC can also be used to find the ground state of general quantum Hamiltonians. The most important requirement is that the ground state is not heated during the process to make it a steady state. This is ensured when the ground state has a finite spectral gap to the excited state and this spectral gap can be resolved by a proper choice of the transition term and the parameters. Consequently, the inverse of the spectral gap determines the basic time scale of the computation. In principle, a finite number of cavity modes can cover the transitions across the entire spectrum of the problem Hamiltonian, where the required number of cavity modes is given roughly by the ratio of the energy range of the problem Hamiltonian to the characteristic transition rate. This means that the number of cavity modes increases polynomially with the system size for gapped Hamiltonians. However, choosing an appropriate transition term and estimating the resulting cooling time depend heavily on the system and are analytically challenging.
Acknowledgements.
This work was supported by the National Research Foundation (NRF) of Korea under Grant No. NRF-2022R1A4A1030660.